# MetaphorShare: A Dynamic Collaborative Repository of Open Metaphor Datasets

### v1

@@ 6 Conclusion and Future Work
In this paper, we have presented a website to unify and share metaphor datasets. The website currently integrates 25 datasets and it enables an easy integration of new resources of different formats by the community. Most of these datasets are in English and we are planning to integrate and reach communities working on other languages in the future as one of the main aims.
One of the main goals of this repository is to join the NLP and metaphor communities to work together, as currently there are common goals but the communities rarely interact. In the future we will focus on an online annotation tool and on integrating a functionality for labelling raw text, optionally based on a set of examples provided by a user. Moreover, the progress recently made by transformer-based language models for metaphor identification opens the path to the creation of helpful personalized tools for semi-automatic labelling of metaphors based on an initial sample of annotations.
@@ Acknowledgments
This research was supported by the Association for Researching and Applying Metaphors through the Building Bridges fund. We thank the British National Corpus for giving us the permission to share fragments of texts of the BNC directly on MetaphorShare. We also thank Eileen Cardillo and all the authors of Gutiérrez et al. (2016) for allowing us to reshare their datasets online for research purposes.
Appendix A Dataset upload
@@
Figure 4 & 5 shows the feedback provided after the automatic validation step.
@@
Figure 2 shows how additional information are displayed on MetaphorShare for the MOH dataset.
@@
Figure 2: Additional information in the MOH dataset integrated in MetaphorShare.
[…]
Figure 3: Screenshot of the datasets information page in the catalog.

### v2

@@ 6 Conclusion and Future Work
In this paper, we have presented a website to label, unify and share metaphor datasets. It enables an easy integration of new resources of different original formats by the community. The 25 currently included datasets are in English and we are planning to integrate and reach communities working on other languages in the future as one of the main aims.
In the future we will focus on integrating a functionality for automatically labelling raw text. The progress recently made by transformer-based language models for metaphor identification opens the path to the creation of helpful personalized tools for automatic or semi-automatic labelling of metaphors based on an initial sample of annotations.
@@ Acknowledgments
This research was supported by the Association for Researching and Applying Metaphors through the Building Bridges fund. We thank the British National Corpus for giving us the permission to share fragments of texts of the BNC directly on MetaphorShare. We also thank Eileen Cardillo, all the authors of Gutiérrez et al. (2016) and of Birke and Sarkar (2006) for allowing us to redistribute their datasets online for research purposes, and to the rest of the authors of the datasets included in the repository for the creation of metaphor datasets with a flexible open license. Without the effort of the metaphor community, this work would not have been possible.
Appendix A Detailed information of datasets upoaded in MetaphorShare and used in our experiments
Table 5 provides detailed information about the datasets, including license, source corpora, label distribution and sentence distribution. Additional links and preprocessing details are listed below :
JANK:
Anomalous sentences are not shown in the table.
MOH:
The license is available at https://saifmohammad.com/WebPages/metaphor.html. The original dataset contains 1639 instances. A few duplicated example sentences caused by orthographic variants of the target word, such as distil/distill have been removed.
NewsMet:
We show the manually labelled sentences (named gold by the authors). The corpus from which the sentences are extracted can be found at https://github.com/several27/FakeNewsCorpus.
TSVET_A:
GUT:
The UKaC is presented in Ferraresi et al. (2008).
PVC:
The original dataset is accessible at https://cogcomp.seas.upenn.edu/page/resource_view/26
MAGPIE:
The PMB corpus from which some sentences are sourced is presented in Abzianidze et al. (2017).
TONG:
Original VUAC sentences and apt (dataset new label) paraphrases are counted in the table.
Appendix B Dataset upload
@@
Figure 3 & 4 show the row by row feedback provided after the automatic validation step when the file is rejected, and the pop up message indicating that a file passed the automatic validation step.
@@
Appendix C Dataset catalog
Figure 2 is a screenshot of the top of the page presenting a dataset in the Catalog section of MetaphorShare. The dataset name, main author and research paper associated to the data release are displayed.
Appendix D Online annotation tool
Figure 5 shows the annotation tool interface with one example sentence. Figure 6 shows how the created records of a dataset are displayed, allowing an easy step by sep construction of a dataset.
@@ Appendix D Online annotation tool
Figure 2: Screenshot of top of the the datasets information page in the catalog section of the website. The English dataset released with Jankowiak (2020) is presented as an example.
@@
[…]
Figure 5: Screenshot of the online annotation tool showing the text input area, tag selection and creation, and coloured resulting tagged text.
[…]
Figure 6: Screenshot showing dataset rows available for tagging or edition, as displayed in the annotation tool.